b) A library contains 60,000 books. 14,000 are on the subject of business, 22,000 are on healthcare and 12,000 on psychology and law. What percentage of the library’s books is on computing, if computing books make up two-thirds of the remainder?


1
Expert's answer
2021-09-30T04:16:15-0400

12/100*12000=8000

8000*100=80000/60000=13.33 percent